find an expression for a viscous forces acting on a Tiny Seed ball of radius R moving in a viscous liquid of velocity n with a constant speed V by the method of dimensional analysis

Grade:11

1 Answers

Rituraj Tiwari
3 years ago
The viscous force F = 6\pi \eta rv

Let, F = k\eta rv

\eta=\dfrac{F}{krv}....(I)

Here, k=6\pi= dimensionless constant

We know that,

F= [MLT^{-2}]

r=[L]

v=[LT^{-1}]

Put the dimension of all element in equation (I)

\eta=\dfrac{F}{krv}

\eta=\dfrac{ [MLT^{-2}]}{[L][LT^{-1}]}

\eta=[ML^{-1}T^{-1}]